OpenStudy (anonymous):

A regular hexagon has a perimeter of 60 m. Find its area. Leave your answer in simplest radical form

OpenStudy (anonymous):

Imagine the hexagon as 6 equilateral triangles stuck together. The outside "boundary" is 60 m, so each of the six triangles has 10-inch sides. The only way I know how to get the altitude(height) of each triangle is to bisect each triangle and use the pythagorean theorem. So, I know the hypotenuse is 10, and the base is half that, so 5 a² + b² = c² a² + 5² = 10² a² + 25 = 100 a² = 100-25 a² = 75 a = 75-² , so the altitude, a is the square root of 75, or 8.66 something. So now we have 12 pieces, each with a=8.66 and b=5 area of triangle = half of the altitude times the base or A=ab/2 8.66``(5)=43.30something divide by 2 and multiply by 12 gets you the total area of the hexagon: 259.80something. Round off and the answer is 260 square meters

OpenStudy (anonymous):

have a good one hope it helped :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yeah np WELCOME TO OPENSTUDY @antc321